OpenStudy (anonymous):

Let C be the curve of intersection of the parabolic cylinder x^2 = 2y and the surface 3z = xy. Find the exact length of C from the origin to the point (6,18,36)

OpenStudy (amistre64):

well, wolfram bites http://www.wolframalpha.com/input/?i=plot+0%3Dxy%2F3+-+z+and+x^2-2y%3D0

OpenStudy (amistre64):

i cant seem to recall parametric to well today

OpenStudy (anonymous):

argh yea i have trouble reparametrizing

OpenStudy (anonymous):

I just dont understand how or what the rationale is for just choosing a random x=t or some other parameter

OpenStudy (amistre64):

\(\frac{x^2}{2} = y\) and the surface 3z = xy \(\frac{x^2}{2} = y\) and the surface z=\(\frac{x^3}{6}\) x y z 0 0 0 1 1/2 1/6 2 2 4/3 3 9/2 9/2 4 8 32/3 5 25/2 125/6 6 18 36 x = t y = t^2/2 z= t^3/6 might define the curve parametrically

OpenStudy (amistre64):

by restating everything in terms of "t"; we can relate it all into one variable

OpenStudy (amistre64):

\[\int f(x,y,z)\sqrt{x'^2+y'^2+z'^2}dt\] \[\int f(x,y,z)\sqrt{1+t^2+\frac{t^4}{4}}dt\] \[\frac{1}{2}\int f(x,y,z)\sqrt{4+4t^2+t^4}dt\] i just cant seem to recall what to do with the f(x,y,z)

OpenStudy (amistre64):

if i take my lead from this: http://tutorial.math.lamar.edu/Classes/CalcIII/VectorArcLength.aspx it says we just need to integrate the tangent vectors mangnitude

OpenStudy (amistre64):

\[\frac{1}{2}\int_{0}^{6}\sqrt{4+4t^2+t^4}dt\] i really do need to brush up on my line integrals i spose